OpenStudy (anonymous):

The volume of a cube is found by using the formula l^3 where l is the side length. If the side length is 4x^3 , what is the volume of the cube? A. 64x^6 B. 12x^9 C. 4x^6 D. 64x^9

OpenStudy (unklerhaukus):

\[V=l^3,\qquad l =4x^3\] What is \(V\) in terms of \(x\)?

OpenStudy (unklerhaukus):

\[V=\big(4x^3\big)^3 = \quad . . .\]

OpenStudy (anonymous):

64x^27

OpenStudy (unklerhaukus):

The index laws you have to use are \[(xy)^m=x^my^m\] and \[(x^m)^n=x^{mn}\]

OpenStudy (anonymous):

64^9 ?

OpenStudy (unklerhaukus):

you didn't type the x , but i know what you mean , and that is right

OpenStudy (anonymous):

thx